Circular No. 8394

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2004 Q2 Donald Edward Machholz reports his visual discovery of a comet (his

tenth) with a 0.15-m f/8 reflector, at 30×, on Aug. 27.467 UT.

2004 UT  α_{2000}  δ_{2000}  Mag. Observer
Aug. 27.511  4^{h}16.^{m}8  −22^{o}20′  11.2 Machholz
27.660  4 16 28.^{s}5 −22 22 50′′  Garradd
27.70611 4 16 32.06 −22 23 01.5  McNaught
27.70738 4 16 32.16 −22 23 01.9  _{”}
27.70867 4 16 32.27 −22 23 02.4  _{”}
27.81640 4 16 40.50 −22 23 36.4  _{”}
27.81766 4 16 40.59 −22 23 36.8  _{”}
27.81894 4 16 40.69 −22 23 37.2  _{”}

D. E. Machholz (Colfax, CA). Coma diameter 2′, moderate degree of condensation (position and physical measurements made with a 0.25-m reflector at 64×). Motion < 1^{o}/day eastward.

G. Garradd (Siding Spring). 0.5-m Uppsala Schmidt telescope + CCD. Coma diameter 1′, with a 3′ tail in p.a. ∼ 320^{o}. Position rough (center of comet uncertain within overexposed coma).

R. H. McNaught and G. Garradd (Siding Spring). 1.0-m f/8 reflector + CCD; 6′ tail in p.a. 245^{o} (10-sec exposure); magnitude of nuclear condensation 14.4.

COMETS C/2004 N1–N6, O1–O3 (SOHO)

Continuation to IAUC 8380 (all Kreutz sungrazers):

Comet  2004 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2004 N1 July 1.300 6^{h}35.^{m}0 +21^{o}46′ C2  XZ 2004-P19
C/2004 N2  1.821 6 35.4 +20 30 C3/2 JS 2004-P19
C/2004 N3  5.488 6 47.5 +19 50 C3/2 JS 2004-P19
C/2004 N4 7.321 6 55.5 +20 25 C2 XZ 2004-P19
C/2004 N5  8.071 6 58.8 +20 18 C3/2 XL 2004-P19
C/2004 N6  9.263 7 02.9 +20 41 C3  XZ 2004-P31
C/2004 O1  16.988 7 24.4 +17 41 C3  XZ 2004-Q11
C/2004 O2  17.638 7 32.6 +18 33 C3  XZ 2004-Q11
C/2004 O3  26.071 8 07.6 +18 15 C3  XZ 2004-Q11
2004 August 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green